### v1

@@ 1 Methods
To perform ideation, the master orchestrator of the system delegates to an idea agent responsible for the formulation of research hypotheses. For the results presented here, we provided seed guidance to the idea agent by specifying the available cognitive tasks and questionnaires, constrained by ethical approval requirements and our domain expertise to ensure validation of each research stage. Such guidance is optional, and if not provided, the agent instead independently targets research fields it considers relevant as the basis for subsequent ideation. The idea agent coordinates several further agents including a review agent which evaluates ideas, a novelty agent which communicates with an archivist agent to perform literature searches via academic databases, and a feasibility agent that assesses methodological constraints and implementation requirements. This multi-agent collaboration aims to ensure that the ideas are scientifically sound, contribute new knowledge, and can be implemented successfully, with ideas passing all checks being constructed into a multifaceted idea framework, which extends that used in previous work38. The finalised ideas are ranked through a multi-stage, multi-model tournament process, ultimately identifying a single ‘best’ idea recommended for further investigation. The full ideation process is summarised in Figure 4.
@@
Upon receiving a validated research idea, the master orchestrator delegates it to a method agent responsible for developing an experimental protocol. Through multiple cycles of evaluation with a review agent, a final consolidated plan is optimised for scientific validity, reliability, and robustness. For studies identified as requiring a priori statistical sample size determination, the method agent engages a power analysis agent instead of relying on arbitrary conventions. Specifically, the power analysis agent independently conducts the necessary calculations to determine a sample size for the research which satisfies scientific standards while meeting practical constraints. To do so, it coordinates the efforts of several specialists: an archivist agent that queries published literature through the search engine tool to establish the essential parameters for power calculations (including anticipated effect sizes, variance estimates, and appropriate statistical tests); a coding agent that writes and executes power analysis code scripts; and a troubleshooting agent that inspects the validity of the outputs and results to provide feedback. Upon completion, the method agent engages a pre-registration agent, which develops a pre-registration report compliant with Open Science Framework (OSF) standards56, detailing hypotheses, independent and dependent variables, sampling procedures, exclusion criteria, analytical approaches, and anticipated outcomes. A review agent then evaluates the proposed research plan, corrects any issues found, and revises the power analysis as necessary. This methodological design framework culminates in a finalised experimental protocol that is designed to meet established methodological standards in the field and provide transparency from the earliest stages of the scientific process.
@@
For experiments requiring physical-world validation, the master orchestrator engages an implementation agent to manage the execution of the experiment within the confines of the specific tools for which ethics approval has been obtained. For the initial experiment, we focussed on online cognitive psychology paradigms, supporting a previously validated Visual Working Memory Task52, Mental Rotation Task53 (MRT), and the Vividness of Visual Imagery Questionnaire-254 (VVIQ2). These were delivered to the agent as a single HTML and JavaScript experiment hosted on Pavlovia (GitLab) with integrated participant information, consent, and a SurveyMonkey questionnaire. The system was also interfaced with Prolific – a high-quality, well-established participant recruitment platform for online experiments109, 110, 111 – which is compatible with web-based Pavlovia experiments.
@@
The master orchestrator passes experimental data onto a dedicated data analysis agent that transforms raw observations into interpretable scientific evidence, prioritising reproducibility, statistical rigour, and explanatory clarity. It manages a multi-stage analytical workflow, collaborating initially with an exclusions agent and archivist agent to establish theoretically informed, literature-backed data cleaning protocols, as well as a validation agent to ensure alignment of the plan with pre-registration commitments. Each analytical stage is subsequently delegated to a coding agent which uses a custom code editing framework based on the file editing functionality of the Aider coding assistant112, to develop the codebase for the step in a structured environment. Each code execution attempt is evaluated by a team of troubleshooting agents, who offer the coding agent diverse analytical perspectives and feedback on implementation challenges. Progress is documented, providing transparent analytical decision paths. Code refinement continues until the troubleshooting agents and independent verification agents reach consensus that the outputs are satisfactory, or when the predefined iteration limit is reached. The data analysis agent subsequently either proceeds to the next analytical steps until the entire data analysis pipeline has been completed, or requests guidance from the researcher if no further progress is being made. The resulting analytical workflow documents detailed interpretations of the results and the full analytical decision process.
@@
The master orchestrator delegates the process of writing a full research report to a manuscript agent, which leverages the skills of specialists to produce a coherent and high-quality writeup. By engaging an archivist agent, the results are situated within the broader scientific literature, establishing connections with existing theoretical frameworks, and constructing a contextual foundation that enhances the explanatory value of the findings. The manuscript agent also develops a logical scaffold for the scientific narrative, delegating aspects to several specialist writing agents who synthesise all the research components – hypotheses, methodological details, results, visualisations, and theoretical implications – into report sections in collaboration with an archivist agent that provides additional relevant literature references as needed. To address the pervasive tendency of LLMs to generate fictitious citations113, 114, 115, the writing agents utilise a validation tool built on doi.org to verify the authenticity of each reference, enabling correction of fabricated or inaccurate references prior to incorporation in the full manuscript. A review agent provides evaluation of the initial writeup version, guided by domain-specific standards, identifying potential improvements in clarity, logical structure, methodological reporting, and theoretical integration. Based on the feedback, the writing agent incorporates targeted refinements, developing a final manuscript to communicate findings.
@@
The master orchestrator delegates final manuscript assembly to a document agent responsible for incorporating all text, figures, tables, and captions into a single publication-ready file. A multi-format implementation allows researchers to specify a preferred format between LaTeX or Microsoft Word. The LaTeX agent constructs its document in a process that builds upon prior work38, with modifications to the structure, layout, formatting, and construction of the document, in addition to building out compatibility, user flexibility and robustness. The Microsoft Word agent utilises a similar process to construct its document, with both formats ultimately also being saved to a final PDF format. Several time-consuming technical aspects are handled by the document agents, including placement of figures and tables, section and caption numbering, encoding of mathematical notation and special characters, reference formatting and typographical standardisation, to produce submission-ready manuscripts that adhere to scientific publishing standards.
@@
Central to the system is frequent evaluation, which functions as a key regulatory mechanism integrated throughout the entire research pipeline. The architecture underlying the review process was derived from that of Paper Wizard – the ‘peer’ review tool – which we tailored to create specialised evaluation protocols for each stage of the scientific process. Feedback allows the system to iteratively refine and adapt its efforts116, 117 until the high standards required for robust scientific inquiry in the field are met. As explicit evaluation and refinement cycles develop higher-quality outputs than single-pass generation73, 83, such intra- and inter-agent feedback mechanisms are fundamental to ensure and maintain methodological rigour and theoretical validity in AI-driven research118. However, it is possible that as the intelligence of LLMs continues to advance, the need for extensive review may eventually diminish.
@@ 2 Results
Comprehensive and relevant literature engagement. Literature engagement was thorough and well-integrated into theoretical framing. Study 1 reviewed the historical development of VWM and mental rotation theory, providing detailed context for the research questions. Study 2 noted the ”reliability paradox” in individual differences research55, demonstrating awareness of contemporary methodological discussions. Study 3 effectively contextualised findings in relation to a recent failed replication of foundational results119.
@@
Theoretical misrepresentations and overstatement. Theoretical models were occasionally misrepresented, including inappropriate extension of a VWM framework120 to unrelated visuospatial tasks in Study 3, and Study 2’s false claim about working memory resource allocation debates. Study 3 claimed a ”novel” imagery-precision link without acknowledging related accuracy research; the statement is misleading given existing literature showing a relationship between imagery strength and VWM accuracy121. Studies 2 and 3 also made interpretative claims unsupported by the data.
@@
The system particularly excelled in rigour and comprehensiveness. Across the manuscripts, there was a consistent emphasis on transparency, robustness checks, and consideration of alternative explanations. Scientists routinely face interpretative challenges when evaluating statistical outputs, particularly given that sufficiently large sample sizes can yield statistical significance for trivially small effects122, 123 where the meaningful magnitude of difference between data distributions is slight. When faced with statistically significant findings but small effect sizes in Study 3, the system demonstrated notable objectivity and sophistication by prioritising practical significance124 over statistical significance alone – a commendable departure from the pervasive p-value fixation that has long plagued human scientific research125, 126, though it potentially risks undervaluing theoretically meaningful discoveries with modest but reliable effects.
@@ 3 Discussion
This study presents an end-to-end AI scientific discovery system that integrates analogues for abstraction, decomposition, metacognition, autonomy, and dynamic memory; and that navigates the complete scientific workflow from hypothesis formulation to manuscript preparation. While domain-specific AI systems have achieved high performance in narrow tasks, such as AlphaFold’s protein structure prediction17, as far as we are aware, this is the first demonstration of an autonomously conducted, end-to-end online experiment with human participants. The results, contained within the three scientific manuscripts that accompany this article, show that AI systems can conduct scientific inquiries with minimal human intervention. While prior systems have primarily operated within computational domains (simulations and modelling), the system’s ability to collect and interpret real-world experimental data from human participants marks a key step toward embodied scientific AI that is capable of testing hypotheses in real-world settings. These findings bridge the gap between in silico AI capabilities and practical scientific application, with important implications for both the future of AI-augmented scientific discovery and fundamental questions about the nature of knowledge creation.
@@
Rigour. Across the three studies, the system tended toward conservative methodological choices, appropriate statistical power, and transparent reporting of limitations – indicating impartiality and cognitive flexibility which can at times challenge human researchers. Importantly, the system’s documentation of all analytical decisions and availability of raw data enables reproducibility, which goes some way toward addressing reproducibility concerns in scientific literature where 70% of researchers have failed to reproduce another scientist’s experiments127, 128, 129. The ability to replicate experimental conditions and analyses at negligible marginal cost could alter how we validate scientific findings, and aligns with the growing movement toward open science practices130.
Sophistication. The system demonstrated the capacity to autonomously construct and execute multi-step data analysis pipelines, employing valid statistical techniques, evaluation of underlying assumptions, and measured interpretation of results. The system here sustained coherent coding and statistical reasoning for longer than eight hours – for context, a $\sim$50 – 55-minute 50% task-completion time horizon has been reported for Claude 3.7 Sonnet on research-engineering tasks131.
@@
Despite its achievements, the system exhibits several limitations that highlight areas for future development. Experimental implementation poses a fundamental bottleneck in the verification of ideas of autonomous science systems132. While the framework is general by design, currently its physical capabilities are domain-constrained to online experiments for which the necessary tools and interfaces are available. Representing an engineering challenge rather than a fundamental limitation, the extension of the system to other scientific fields requires the development and incorporation of new and existing domain-specific toolsets, which are readily accommodated by the modular and generalist core system architecture.
@@
Furthermore, while the system was capable of end-to-end research, it was not infallible and can thus benefit from human refinement and oversight. In particular, data visualisations occasionally include graphical imperfections which persist despite inspection and verification mechanisms – such as overlapping axis labels, misaligned label positioning, or suboptimal axis bounds. Though predominantly aesthetic rather than substantively inaccurate, these minor discrepancies highlight the intrinsic challenges in automated computer vision tasks that human visual perception resolves effortlessly133, 134. Such artifacts remain difficult for the system to detect autonomously, yet can be quickly and easily rectified with minimal human intervention.
@@
In addition, rigorous scientific inquiry involves lengthy, complex, multi-stage processes, requiring the system to handle extremely long chains of thought – thousands of reasoning steps and conversational exchanges extending over 12 hours. The dynamic memory system and cognitive offloading mechanisms were central to maintaining conceptual continuity across research stages, effectively combining working memory with strategic external resource utilisation. The system’s ability to selectively filter relevant information while preserving focused knowledge representations of the necessary context enabled it to navigate the entire scientific workflow without conceptual drift. Interestingly, many reasoning models generally degrade in their performance135, 136 over extremely long chains, losing focus and coherence across iterations.
@@
Sensitivity to early-stage accuracy also emerged as a challenge. Poor question formulation or any conceptual errors introduced during hypothesis generation and methodological design propagate downstream – persisting through multiple verification cycles – at a detriment to research outcomes. This potential single‑point failure mode likely reflects the “anchoring” bias characterised in LLMs137, whereby the first logical claims encountered by the model are weighted as ground truth and only weakly revised later, disproportionately shaping subsequent judgements. Consequently, once a misconception enters the chain of thought, later processes tend to build upon it rather than correcting it138. Explicit error-checking steps only seldom succeed in reversing the trajectory once a false premise has been internalised. Interestingly, humans also exhibit this cognitive bias139, 140, 141, where initial information acts as a reference point that substantially affects how subsequent information is assessed, particularly under uncertainty. This phenomenon emphasises the importance of robust verification protocols during the earliest conceptual stages of automated scientific investigation and highlights a crucial advantage of human-AI collaborative scientific workflows where human expertise can intervene most effectively at points of highest conceptual leverage.
@@
Models with real-world experimental capabilities represent a potential pathway toward more advanced AI. The framework we have developed embodies the fundamental cycle of knowledge creation: executing movements to explore the world, reflection on outcomes, understanding relationships with prior knowledge, generating insights from these interactions, and developing foresight for future predictions142. This process mirrors cognitive development in children, who learn primarily through physical manipulation of objects and progressive refinement of their understanding105, 143. While current LLMs excel at pattern recognition within training data, they remain limited by their inability to autonomously expand beyond those boundaries. Here the system coupled internal representations – the exploration of latent connections between concepts – with external measurement through experimentation. This creates a virtuous cycle where hypotheses conceived in the model’s rich latent space can be validated against reality, with results feeding back to refine its conceptual framework. This suggests a fundamentally different approach to advancing AI capabilities – one rooted in the power of scientific investigation to build increasingly accurate models of the world. Toward this end, we use the intermediary milestone of ‘Artificial General Science’ (AGS) to denote autonomous systems capable of independently driving scientific inquiry across all domains – generating hypotheses, orchestrating experiments, and iteratively refining knowledge through empirical evidence. In a recursive loop of discovery and learning, the scope of understanding and knowledge of the system could be continually expanded beyond trained-on human knowledge.
@@
Autonomous scientific systems capable of designing and executing rigorous methodological plans will likely reshape the role of human scientists144. While capable of independent operation with minimal human intervention, the balance between autonomy and human collaboration offers distinct advantages depending on the research goal. Currently, humans provide most value to these systems in creative problem formulation, conceptual innovation, and ethical oversight – though the focal points for human contribution may shift as the technology advances. These systems can also address limitations of contemporary research, particularly the increasingly narrow specialisation that often constrains scientists to their areas of expertise145, impeding progress on complex interdisciplinary problems. By bridging disciplinary boundaries, domain-agnostic frameworks facilitate integrative research, empowering scientists to explore theoretically adjacent areas where they may lack training but possess valuable conceptual insights. Researchers can delegate any aspect(s) of the scientific workflow to the system as needed, enhancing both efficiency and innovation through collaboration.
@@
The accelerated pace of research enabled by autonomous scientific systems presents both opportunities and challenges. While rapid knowledge generation could expedite solutions to pressing global issues from climate change to disease, inadequate consideration of downstream effects risks unintended consequences. Of particular concern is the potential for producing research outputs at a volume that overwhelms human researchers, necessitating innovative approaches to distil and communicate findings effectively within human cognitive capacity constraints. Notably, autonomous scientific systems are likely to generate a high proportion of null findings, which, while typically remaining unpublished in traditional research146, offer under-utilised value. By documenting non-significant outcomes, these results may mitigate publication bias146, as well as characterise the ”null space” of research fields – highlighting where relationships are absent and conversely where they may exist. This dual benefit reduces wasted resources on redundant experiments and can guide researchers toward better-informed hypotheses for promising investigations. The environmental impact of running multiple LLMs over sustained periods also warrants consideration, as the material energy and carbon footprint of such systems is non-trivial, and yet to be quantified relative to the net sustainability of human labour. Balancing the dynamics of speed, volume, environmental impact, and meaningful human oversight will be important as autonomous discovery systems become more prevalent in scientific workflows.

### v2

@@ 1 Methods
To perform ideation, the master orchestrator of the system delegates to an idea agent responsible for the formulation of research hypotheses. For the results presented here, we provided seed guidance to the idea agent by specifying the available cognitive tasks and questionnaires, constrained by ethical approval requirements and our domain expertise to ensure validation of each research stage. Such guidance is optional, and if not provided, the agent instead independently targets research fields it considers relevant as the basis for subsequent ideation. The idea agent coordinates several further agents including a review agent which evaluates ideas, a novelty agent which communicates with an archivist agent to perform literature searches via academic databases, and a feasibility agent that assesses methodological constraints and implementation requirements. This multi-agent collaboration aims to ensure that the ideas are scientifically sound, contribute new knowledge, and can be implemented successfully, with ideas passing all checks being constructed into a multifaceted idea framework, which extends that used in previous worklu_ai_2024. The finalised ideas are ranked through a multi-stage, multi-model tournament process, ultimately identifying a single ‘best’ idea recommended for further investigation. The full ideation process is summarised in Figure 4.
@@
Upon receiving a validated research idea, the master orchestrator delegates it to a method agent responsible for developing an experimental protocol. Through multiple cycles of evaluation with a review agent, a final consolidated plan is optimised for scientific validity, reliability, and robustness. For studies identified as requiring a priori statistical sample size determination, the method agent engages a power analysis agent instead of relying on arbitrary conventions. Specifically, the power analysis agent independently conducts the necessary calculations to determine a sample size for the research which satisfies scientific standards while meeting practical constraints. To do so, it coordinates the efforts of several specialists: an archivist agent that queries published literature through the search engine tool to establish the essential parameters for power calculations (including anticipated effect sizes, variance estimates, and appropriate statistical tests); a coding agent that writes and executes power analysis code scripts; and a troubleshooting agent that inspects the validity of the outputs and results to provide feedback. Upon completion, the method agent engages a pre-registration agent, which develops a pre-registration report compliant with Open Science Framework (OSF) standardsfoster_open_2017, detailing hypotheses, independent and dependent variables, sampling procedures, exclusion criteria, analytical approaches, and anticipated outcomes. A review agent then evaluates the proposed research plan, corrects any issues found, and revises the power analysis as necessary. This methodological design framework culminates in a finalised experimental protocol that is designed to meet established methodological standards in the field and provide transparency from the earliest stages of the scientific process.
@@
For experiments requiring physical-world validation, the master orchestrator engages an implementation agent to manage the execution of the experiment within the confines of the specific tools for which ethics approval has been obtained. For the initial experiment, we focussed on online cognitive psychology paradigms, supporting a previously validated Visual Working Memory Taskzhang_discrete_2008, Mental Rotation Taskshepard_mental_1971 (MRT), and the Vividness of Visual Imagery Questionnaire-2marks_new_1995 (VVIQ2). These were delivered to the agent as a single HTML and JavaScript experiment hosted on Pavlovia (GitLab) with integrated participant information, consent, and a SurveyMonkey questionnaire. The system was also interfaced with Prolific – a high-quality, well-established participant recruitment platform for online experimentspalan_prolificacsubject_2018, peer_beyond_2017, peer_data_2021 – which is compatible with web-based Pavlovia experiments.
@@
The master orchestrator passes experimental data onto a dedicated data analysis agent that transforms raw observations into interpretable scientific evidence, prioritising reproducibility, statistical rigour, and explanatory clarity. It manages a multi-stage analytical workflow, collaborating initially with an exclusions agent and archivist agent to establish theoretically informed, literature-backed data cleaning protocols, as well as a validation agent to ensure alignment of the plan with pre-registration commitments. Each analytical stage is subsequently delegated to a coding agent which uses a custom code editing framework based on the file editing functionality of the Aider coding assistantgauthier_aider_2025, to develop the codebase for the step in a structured environment. Each code execution attempt is evaluated by a team of troubleshooting agents, who offer the coding agent diverse analytical perspectives and feedback on implementation challenges. Progress is documented, providing transparent analytical decision paths. Code refinement continues until the troubleshooting agents and independent verification agents reach consensus that the outputs are satisfactory, or when the predefined iteration limit is reached. The data analysis agent subsequently either proceeds to the next analytical steps until the entire data analysis pipeline has been completed, or requests guidance from the researcher if no further progress is being made. The resulting analytical workflow documents detailed interpretations of the results and the full analytical decision process.
@@
The master orchestrator delegates the process of writing a full research report to a manuscript agent, which leverages the skills of specialists to produce a coherent and high-quality writeup. By engaging an archivist agent, the results are situated within the broader scientific literature, establishing connections with existing theoretical frameworks, and constructing a contextual foundation that enhances the explanatory value of the findings. The manuscript agent also develops a logical scaffold for the scientific narrative, delegating aspects to several specialist writing agents who synthesise all the research components – hypotheses, methodological details, results, visualisations, and theoretical implications – into report sections in collaboration with an archivist agent that provides additional relevant literature references as needed. To address the pervasive tendency of LLMs to generate fictitious citationsbhattacharyya_high_2023, walters_fabrication_2023, wu_automated_2025, the writing agents utilise a validation tool built on doi.org to verify the authenticity of each reference, enabling correction of fabricated or inaccurate references prior to incorporation in the full manuscript. A review agent provides evaluation of the initial writeup version, guided by domain-specific standards, identifying potential improvements in clarity, logical structure, methodological reporting, and theoretical integration. Based on the feedback, the writing agent incorporates targeted refinements, developing a final manuscript to communicate findings.
@@
The master orchestrator delegates final manuscript assembly to a document agent responsible for incorporating all text, figures, tables, and captions into a single publication-ready file. A multi-format implementation allows researchers to specify a preferred format between LaTeX or Microsoft Word. The LaTeX agent constructs its document in a process that builds upon prior worklu_ai_2024, with modifications to the structure, layout, formatting, and construction of the document, in addition to building out compatibility, user flexibility and robustness. The Microsoft Word agent utilises a similar process to construct its document, with both formats ultimately also being saved to a final PDF format. Several time-consuming technical aspects are handled by the document agents, including placement of figures and tables, section and caption numbering, encoding of mathematical notation and special characters, reference formatting and typographical standardisation, to produce submission-ready manuscripts that adhere to scientific publishing standards.
@@
Central to the system is frequent evaluation, which functions as a key regulatory mechanism integrated throughout the entire research pipeline. The architecture underlying the review process was derived from that of Paper Wizard – the ‘peer’ review tool – which we tailored to create specialised evaluation protocols for each stage of the scientific process. Feedback allows the system to iteratively refine and adapt its effortsashby_introduction_1956, wiener_cybernetics_1948 until the high standards required for robust scientific inquiry in the field are met. As explicit evaluation and refinement cycles develop higher-quality outputs than single-pass generationwei_chain--thought_2022, madaan_self-refine_2023, such intra- and inter-agent feedback mechanisms are fundamental to ensure and maintain methodological rigour and theoretical validity in AI-driven researchkon_curie_2025. However, it is possible that as the intelligence of LLMs continues to advance, the need for extensive review may eventually diminish.
@@ 2 Results
Comprehensive and relevant literature engagement. Literature engagement was thorough and well-integrated into theoretical framing. Study 1 reviewed the historical development of VWM and mental rotation theory, providing detailed context for the research questions. Study 2 noted the ”reliability paradox” in individual differences researchhedge_reliability_2018, demonstrating awareness of contemporary methodological discussions. Study 3 effectively contextualised findings in relation to a recent failed replication of foundational resultsebert_visual_2025.
@@
Theoretical misrepresentations and overstatement. Theoretical models were occasionally misrepresented, including inappropriate extension of a VWM frameworkbays_precision_2009 to unrelated visuospatial tasks in Study 3, and Study 2’s false claim about working memory resource allocation debates. Study 3 claimed a ”novel” imagery-precision link without acknowledging related accuracy research; the statement is misleading given existing literature showing a relationship between imagery strength and VWM accuracyjacobs_visual_2018. Studies 2 and 3 also made interpretative claims unsupported by the data.
@@
The system particularly excelled in rigour and comprehensiveness. Across the manuscripts, there was a consistent emphasis on transparency, robustness checks, and consideration of alternative explanations. Scientists routinely face interpretative challenges when evaluating statistical outputs, particularly given that sufficiently large sample sizes can yield statistical significance for trivially small effectswasserstein_moving_2019, wasserstein_asa_2016 where the meaningful magnitude of difference between data distributions is slight. When faced with statistically significant findings but small effect sizes in Study 3, the system demonstrated notable objectivity and sophistication by prioritising practical significancekirk_practical_1996 over statistical significance alone – a commendable departure from the pervasive p-value fixation that has long plagued human scientific researchcohen_earth_1994, sullivan_using_2012, though it potentially risks undervaluing theoretically meaningful discoveries with modest but reliable effects.
@@ 3 Discussion
This study presents an end-to-end AI scientific discovery system that integrates analogues for abstraction, decomposition, metacognition, autonomy, and dynamic memory; and that navigates the complete scientific workflow from hypothesis formulation to manuscript preparation. While domain-specific AI systems have achieved high performance in narrow tasks, such as AlphaFold’s protein structure predictionjumper_highly_2021, as far as we are aware, this is the first demonstration of an autonomously conducted, end-to-end online experiment with human participants. The results, contained within the three scientific manuscripts that accompany this article, show that AI systems can conduct scientific inquiries with minimal human intervention. While prior systems have primarily operated within computational domains (simulations and modelling), the system’s ability to collect and interpret real-world experimental data from human participants marks a key step toward embodied scientific AI that is capable of testing hypotheses in real-world settings. These findings bridge the gap between in silico AI capabilities and practical scientific application, with important implications for both the future of AI-augmented scientific discovery and fundamental questions about the nature of knowledge creation.
@@
Rigour. Across the three studies, the system tended toward conservative methodological choices, appropriate statistical power, and transparent reporting of limitations – indicating impartiality and cognitive flexibility which can at times challenge human researchers. Importantly, the system’s documentation of all analytical decisions and availability of raw data enables reproducibility, which goes some way toward addressing reproducibility concerns in scientific literature where 70% of researchers have failed to reproduce another scientist’s experimentsbaker_1500_2016, camerer_evaluating_2018, open_science_collaboration_estimating_2015. The ability to replicate experimental conditions and analyses at negligible marginal cost could alter how we validate scientific findings, and aligns with the growing movement toward open science practicesgong_open_2022.
Sophistication. The system demonstrated the capacity to autonomously construct and execute multi-step data analysis pipelines, employing valid statistical techniques, evaluation of underlying assumptions, and measured interpretation of results. The system here sustained coherent coding and statistical reasoning for longer than eight hours – for context, a $\sim$50 – 55-minute 50% task-completion time horizon has been reported for Claude 3.7 Sonnet on research-engineering taskskwa_measuring_2025.
@@
Despite its achievements, the system exhibits several limitations that highlight areas for future development. Experimental implementation poses a fundamental bottleneck in the verification of ideas of autonomous science systemszhu_ai_2025. While the framework is general by design, currently its physical capabilities are domain-constrained to online experiments for which the necessary tools and interfaces are available. Representing an engineering challenge rather than a fundamental limitation, the extension of the system to other scientific fields requires the development and incorporation of new and existing domain-specific toolsets, which are readily accommodated by the modular and generalist core system architecture.
@@
Furthermore, while the system was capable of end-to-end research, it was not infallible and can thus benefit from human refinement and oversight. In particular, data visualisations occasionally include graphical imperfections which persist despite inspection and verification mechanisms – such as overlapping axis labels, misaligned label positioning, or suboptimal axis bounds. Though predominantly aesthetic rather than substantively inaccurate, these minor discrepancies highlight the intrinsic challenges in automated computer vision tasks that human visual perception resolves effortlesslydicarlo_how_2012, geirhos_shortcut_2020. Such artifacts remain difficult for the system to detect autonomously, yet can be quickly and easily rectified with minimal human intervention.
@@
In addition, rigorous scientific inquiry involves lengthy, complex, multi-stage processes, requiring the system to handle extremely long chains of thought – thousands of reasoning steps and conversational exchanges extending over 12 hours. The dynamic memory system and cognitive offloading mechanisms were central to maintaining conceptual continuity across research stages, effectively combining working memory with strategic external resource utilisation. The system’s ability to selectively filter relevant information while preserving focused knowledge representations of the necessary context enabled it to navigate the entire scientific workflow without conceptual drift. Interestingly, many reasoning models generally degrade in their performanceballon_relationship_2025, lin_zebralogic_2025 over extremely long chains, losing focus and coherence across iterations.
@@
Sensitivity to early-stage accuracy also emerged as a challenge. Poor question formulation or any conceptual errors introduced during hypothesis generation and methodological design propagate downstream – persisting through multiple verification cycles – at a detriment to research outcomes. This potential single‑point failure mode likely reflects the “anchoring” bias characterised in LLMslou_anchoring_2024, whereby the first logical claims encountered by the model are weighted as ground truth and only weakly revised later, disproportionately shaping subsequent judgements. Consequently, once a misconception enters the chain of thought, later processes tend to build upon it rather than correcting itli_beyond_2025. Explicit error-checking steps only seldom succeed in reversing the trajectory once a false premise has been internalised. Interestingly, humans also exhibit this cognitive biasfurnham_literature_2011, strack_explaining_1997, tversky_judgment_1974, where initial information acts as a reference point that substantially affects how subsequent information is assessed, particularly under uncertainty. This phenomenon emphasises the importance of robust verification protocols during the earliest conceptual stages of automated scientific investigation and highlights a crucial advantage of human-AI collaborative scientific workflows where human expertise can intervene most effectively at points of highest conceptual leverage.
@@
Models with real-world experimental capabilities represent a potential pathway toward more advanced AI. The framework we have developed embodies the fundamental cycle of knowledge creation: executing movements to explore the world, reflection on outcomes, understanding relationships with prior knowledge, generating insights from these interactions, and developing foresight for future predictionskolb_experiential_1984. This process mirrors cognitive development in children, who learn primarily through physical manipulation of objects and progressive refinement of their understandingpiaget_origins_1952, gopnik_theory_2004. While current LLMs excel at pattern recognition within training data, they remain limited by their inability to autonomously expand beyond those boundaries. Here the system coupled internal representations – the exploration of latent connections between concepts – with external measurement through experimentation. This creates a virtuous cycle where hypotheses conceived in the model’s rich latent space can be validated against reality, with results feeding back to refine its conceptual framework. This suggests a fundamentally different approach to advancing AI capabilities – one rooted in the power of scientific investigation to build increasingly accurate models of the world. Toward this end, we use the intermediary milestone of ‘Artificial General Science’ (AGS) to denote autonomous systems capable of independently driving scientific inquiry across all domains – generating hypotheses, orchestrating experiments, and iteratively refining knowledge through empirical evidence. In a recursive loop of discovery and learning, the scope of understanding and knowledge of the system could be continually expanded beyond trained-on human knowledge.
@@
Autonomous scientific systems capable of designing and executing rigorous methodological plans will likely reshape the role of human scientistsking_automation_2009. While capable of independent operation with minimal human intervention, the balance between autonomy and human collaboration offers distinct advantages depending on the research goal. Currently, humans provide most value to these systems in creative problem formulation, conceptual innovation, and ethical oversight – though the focal points for human contribution may shift as the technology advances. These systems can also address limitations of contemporary research, particularly the increasingly narrow specialisation that often constrains scientists to their areas of expertisejones_burden_2009, impeding progress on complex interdisciplinary problems. By bridging disciplinary boundaries, domain-agnostic frameworks facilitate integrative research, empowering scientists to explore theoretically adjacent areas where they may lack training but possess valuable conceptual insights. Researchers can delegate any aspect(s) of the scientific workflow to the system as needed, enhancing both efficiency and innovation through collaboration.
@@
The accelerated pace of research enabled by autonomous scientific systems presents both opportunities and challenges. While rapid knowledge generation could expedite solutions to pressing global issues from climate change to disease, inadequate consideration of downstream effects risks unintended consequences. Of particular concern is the potential for producing research outputs at a volume that overwhelms human researchers, necessitating innovative approaches to distil and communicate findings effectively within human cognitive capacity constraints. Notably, autonomous scientific systems are likely to generate a high proportion of null findings, which, while typically remaining unpublished in traditional researchrosenthal_file_1979, offer under-utilised value. By documenting non-significant outcomes, these results may mitigate publication biasrosenthal_file_1979, as well as characterise the ”null space” of research fields – highlighting where relationships are absent and conversely where they may exist. This dual benefit reduces wasted resources on redundant experiments and can guide researchers toward better-informed hypotheses for promising investigations. The environmental impact of running multiple LLMs over sustained periods also warrants consideration, as the material energy and carbon footprint of such systems is non-trivial, and yet to be quantified relative to the net sustainability of human labour. Balancing the dynamics of speed, volume, environmental impact, and meaningful human oversight will be important as autonomous discovery systems become more prevalent in scientific workflows.